OpenStudy (anonymous):

solve by completing the square x^2-8x+32=0

OpenStudy (anonymous):

use x^2-2xy+y^2=(x+y)^2, now just find out what is y in this !

OpenStudy (anonymous):

(x-Y)^2 i meant :D

OpenStudy (anonymous):

(x-4)^2 - 16 + 32 = 0 (x-4)^2 = -16 take sqrt both sides

OpenStudy (anonymous):

ok but when i get 4+or-4i how do I determine if it is just +4i or if it is +or- 4i

OpenStudy (anonymous):

x-4 = 4i or - 4i x = 4 + 4i or 4- 4i done